# Saliency Guided Image Warping for Unsupervised Domain Adaptation

### v2

@@ 3.3 Warping for Unsupervised Domain Adaptation
We consider unsupervised domain adaptation, where the source domain $\mathcal{X}_{a}$ includes labeled data, i.e., $X_{a}=\{\ (x_{a},y_{a})\in\mathcal{X}_{a}\ \}$, whereas the target domain $\mathcal{X}_{b}$ consists only of unlabeled images, i.e., $X_{b}=\{\ x_{b}\in\mathcal{X}_{b}\ \}$.
Domain Adaption Framework: The standard framework follows a student-teacher self-training approach. It begins with a supervised pre-training stage where a model $\hat{y}_{a}=F_{\theta}(x_{a})$ is trained using supervised loss on the source domain, $\mathcal{L}_{sup}(y_{a},F_{\theta}(x_{a})).$ Next, in the unsupervised adaptation stage, the model is adapted to the target domain $\mathcal{X}_{b}$ using dataset $X_{b}$ through an additional unsupervised loss. Generally, a student $F_{s,\phi}$ and teacher $F_{t,\kappa}$ (initialized using $F_{\theta}$) are trained jointly through self-training with psuedo-labels [17, 13]. Following [13, 14, 15, 17], we update the teacher parameters, as DINO [3] and Mean Teachers [34] have shown that this approach leads to better performance compared to traditional distillation. While training the student, a supervised loss is used to maintain performance on the source domain $\mathcal{X}_{a}$ in a supervised branch, $\mathcal{L}_{sup}(y_{a},F_{s,\phi}(x_{a}))+\lambda\mathcal{L}_{unsup}(F_{t,\kappa}(x_{b}),F_{s,\phi}(x_{b}))$
@@
Warping and Unwarping in the Domain Adaptation Framework: Our warping mechanism $W$ operates only on source domain images while training, i.e. it is employed in the supervised pre-training stage and in the student model during the unsupervised stage when source images are used. Consider the recognition model $F$ as a combination of backbone $B$ and prediction heads $H$, i.e. $F(x)=H(B(x))$. We warp the image and then unwarp the feature maps from the backbone $B$ to compose the final output,
@@
We have experimented with warping target images (during training and inference) using Static [35, 36] and Geometric [9] prior saliency guidances, which do not require ground truth labels. However, this approach reduced performance (See Table 8). Finally, we do not warp during test time because it introduces additional latency and our ablation study (Table 8) shows that it does not improve performance after adaptation.
Base Algorithm and Losses: We use DAFormer [13] and 2PCNet [17] as our base segmentation and detection adaption algorithms, respectively. We directly adopt their losses for $L_{sup}$ and $L_{unsup}$ without modification. Our method is task-agnostic (segmentation or detection) and makes no additional assumptions about the setup. Therefore, other algorithms [14, 15, 18, 6, 49, 42] are equally applicable.
Inference: We employ the teacher model $F_{t}$ during inference. Our method does not require image warping or feature unwarping during testing. This contrast with image warping techniques proposed in the supervised setting [35, 9, 36], which incur overheads from test-time warping and require additional information (like vanishing points [9]) to generate saliency maps for warping at test time.
Supervised Setting: Our training method can also be applied in supervised learning scenario by following the supervised pre-training steps to train a model $F^{\prime}$. That is, omitting the unsupervised training stage completely and only performing supervised pre-training stage – then removing the warping and unwarping layers completely from the trained model. This is unlike prior work [24, 35, 9, 36] which also advocates warping at test time. With our warping on the source domain, model performance improves on the target domain $\mathcal{X}_{b}$ even without adaptation (See Table 5 and 6). For Source Domain Results, see Appendix C.
@@
4 Experimental Methodology
We implement our saliency warping technique alongside a base domain adaptation algorithm, ensuring fair comparisons by maintaining the same model, datasets, training schedules, hyperparameters, and random seeds.
Semantic Segmentation: We follow DAFormer [13] by employing the same SegFormer [45] head and MiT-B5 [45] backbone, adhereing to their training hyperparameters, protocols, and seed for fair comparison. While Sim2Real Gap result presented in DAFormer [13] are not our focus, relevant results and discussion can be found in Appendix A.
Datasets: We use BDD100K [48], Cityscapes [5], DENSE [2], ACDC [30] and DarkZurich [28] (Appendix F).
@@ 5.1 Domain Adaptive Object Detection
Tables 1, 2,  3 and 4 show improvements over the state-of-the-art 2PCNet [17] across lighting, weather and geographic adaptations. Moreover, our instance-level saliency guidance consistently outperforms approximate saliency guidance [35, 9]. Note that 2PCNet [17] focused solely on Day $\rightarrow$ Night adaptation with its NightAug augmentation, which was only used in the BDD100K Day $\rightarrow$ Night experiment (Table 3) and omitted for others.
Figure 5 shows examples where 2PCNet [17] mistakes parts of the road for cars and street lights for traffic lights (For other domain adaptations, see Appendix E). In contrast, our method accurately distinguish these foreground objects and background element in challenging lighting and weather conditions.
@@ 5.2 Domain Adaptive Semantic Segmentation
To demonstrate the task-agnostic nature of our method, we integrate it into domain adaptive semantic segmentation using the popular DAFormer [13] algorithm. We observe significant improvement on both Val and Test Splits of the ACDC dataset (Table 5). Our instance-level image saliency guidance outperforms Static Prior [35] and Geometric Prior [9]. Being agnostic to the DA algorithm, our method also boosts DAFormer extensions like HRDA [14] and MIC [15] (See Appendix B). These extensions add significant training memory (+5GB) and inference latency (+800 ms on RTX 4090 GPU), whereas our approach requires negligible additional training memory (+0.04 GB) and adds no additional inference latency.
Table 5 also demonstrates our method’s superiority across scenarios, both with and without unsupervised domain adaptation, suggesting the learnt backbone features generalize better to unseen domains. While we briefly discuss some qualitative results using GradCam [31] visualizations, detailed visual diagnosis of the learned backbones is available in Appendix E. Additionally, Figure 6 illustrates the superior performance of our method in identifying both background (e.g., sky, sidewalks, buildings) and objects (e.g., traffic lights, traffic signs).
@@ 6 Limitations and Conclusion
In this work, we demonstrate that saliency guided in-place image warping improves Unsupervised Domain Adaptation (UDA). Unlike previous approaches that indiscriminately process objects and backgrounds, our instance-level saliency guidance oversamples object regions and undersamples background regions. This improves learned backbone features, which in turn helps domain adaptation. Extensive experiments validate the effectiveness of our warping approach and instance-level saliency guidance. Notably, our approach is efficient, with minimal training overhead and no additional inference latency, and is agnostic to the adaptation task, target domain, saliency guidance, model architecture, and domain adaptation algorithm.
One limitation of our method is its reduced effectiveness in densely populated scenes (e.g., Times Square) where there is limited background to reduce for enlarging foreground objects. Our in-place warping strategy is less effective when background is less than one-fourth of the image. Besides, our approach is less effective on synthetic datasets like Synthia [27], captured from various viewpoints. However, it proves effective on other synthetic datasets like GTA5 [26], which feature realistic city driving scenes.
@@ Appendix A Synthetic to Real Domain Adaptation
Table 9: Synthetic to Real Domain Adaptation: GTA $\rightarrow$ Cityscapes Semantic Segmentation. Tested on Cityscapes Val. We were unable to reproduce DAFormer’s [13] results averaged over 3 random seeds. Therefore, we present results with seed=0 for both DAFormer [13] and our method (DAFormer with instance-level saliency guidance). Our method shows improvement in many individual categories for GTA. Although our focus is on real datasets rather than synthetic datasets, GTA closely resembles natural driving scenes due to its realistic traffic simulation and vehicle-perspective imagery (See Figure 7). These realistic object-background appearances with background dominance contribute to the effectiveness of our method.
@@
Table 10: Synthetic to Real Domain Adaptation: Synthia $\rightarrow$ Cityscapes Semantic Segmentation. Tested on Cityscapes Val. We were unable to reproduce DAFormer’s [13] results averaged over three random seeds. Therefore, we present results with seed=0 for both DAFormer [13] and our method (DAFormer with instance-level saliency guidance). Our improvement on Synthia $\rightarrow$ Cityscapes is marginal, which we attribute to the fact that Synthia does not exhibit appearance akin to natural driving scenes (See Figure 7). These unrealistic object-background appearances make our method ineffective.
@@ Appendix B Comparison with other Domain Adaptation Strategies
Table 11: Comparison with other Domain Adaptation Strategies: Computational Efficiency. We use a single RTX 4090 GPU with a batch size of 1, maintaining the same training and inference image sizes as specified in DAFormer [13]. Our saliency-guided image warping imposes minimal memory overhead during training. Additionally, since we do not perform warping during test time, our method incurs no inference latency overhead.
Method Training Memory Inference Time
DAFormer 16.41 GB 195.0 ms
DAFormer + Ours + 0.04 GB + 0 ms
DAFormer + HRDA + 5.11 GB + 801.2 ms
@@ B.2 Computational Efficiency
Table 11 shows the computational efficiency comparison using the DAFormer image scale. We observe that incorporating HRDA into DAFormer significantly increases training and inference computational costs due to HRDA’s train-time HR-cropping and test-time sliding window inference. In contrast, our inference costs remain the same as those of DAFormer [13], as we do not warp during test time. Our training memory consumption are only slightly higher, due to the lightweight design of our warping-related modules. Additionally, our saliency-guided image warping is orthogonal to the aforementioned training strategies and can be utilized together with them.
@@ B.3 Evaluation Methodology
Based on the above discussion, we apply HRDA’s and MIC’s strategies to DAFormer. However, for a fair comparison, we perform training and inference following DAFormer’s [13] training/testing image scales and evaluation paradigm. An alternative comparison would involve training and evaluating on full-scale images to match HRDA, which deviates from original DAFormer [13] training and evaluation that uses half-scaled images. Unfortunately, we lack access to a GPU with high memory capacity, such as the NVIDIA Tesla A40 with 48 GB or Tesla A100 with 80 GB of memory.
@@ Appendix E Additional Analysis
Appendix F Additional Experimental Details
The following are the datasets used in our experiments.

### v3

@@ 3.4 Warping for Unsupervised Domain Adaptation
We first discuss the standard domain adaptation framework, followed by our modifications to incorporate warping and unwarping into this framework.
Domain Adaptation Framework. The standard framework follows a student-teacher self-training approach. It begins with a supervised pre-training stage where a model $\hat{y}_{a}=F_{\theta}(x_{a})$ is trained using supervised loss on the source domain, $\mathcal{L}_{sup}(y_{a},F_{\theta}(x_{a})).$ Next, in the unsupervised adaptation stage, the model is adapted to the target domain $\mathcal{X}_{b}$ using dataset $X_{b}$ through an additional unsupervised loss. Generally, a student $F_{s,\phi}$ and teacher $F_{t,\kappa}$ (initialized using $F_{\theta}$) are trained jointly through self-training with psuedo-labels [19, 15]. The main insight of self-training is that the student and teacher branches are trained with different image augmentations or views, and their predicted labels should match each other. Following [15, 16, 17, 19], we update the teacher parameters, as DINO [4] and Mean Teachers [37] have shown that this approach leads to better performance compared to traditional distillation. While training the student, a supervised loss is used to maintain performance on the source domain $\mathcal{X}_{a}$ in a supervised branch – thus loss is $\mathcal{L}_{sup}(y_{a},F_{s,\phi}(x_{a}))+\lambda\mathcal{L}_{unsup}(F_{t,\kappa}(x_{b}),F_{s,\phi}(x_{b}))$.
@@
Warping and Unwarping for Domain Adaptation. Our warping mechanism $W$ operates only on source domain images while training, i.e. it is employed in the supervised pre-training stage and in the student model during the unsupervised stage when source images are used. Consider the recognition model $F$ as a combination of backbone $B$ and prediction heads $H$, i.e. $F(x)=H(B(x))$. We warp the image and then unwarp the feature maps from the backbone $B$ to compose the final output,
@@
Instance-level saliency requires ground truth labels, making it unfeasible for warping target images during UDA. However, we experimented by warping target images (during training and testing) using Static [38, 39] and Geometric [10] saliency guidance, which do not require ground truth. However, this degrades performance (See Table 8). We hypothesize that warping both source and target images hinders the adaptability of the student model. Therefore, We warp only source domain images, while target domain images remain warped. This encourages the student model to learn to predict in both the warped and unwarped image spaces, as it is exposed to both types of views. In contrast, the teacher model only observes the unwarped views. The disparity between the teacher’s and the student’s view can be regarded as a form of data augmentation. We leave the analysis of the effect of this teacher-student disparity, which has also been explored in works like [4, 11, 15, 17, 16, 19], for future work. Finally, we do not warp during test time because it introduces additional latency and our ablation study (Table 8) shows that it does not improve performance.
@@
Base Algorithm and Losses. We use DAFormer [15] and 2PCNet [19] as our base segmentation and detection adaption algorithms, respectively. We directly adopt their losses for $L_{sup}$ and $L_{unsup}$ without modification. Our method is task-agnostic (segmentation or detection). Therefore, other algorithms [16, 17, 21, 7, 53, 46] are equally applicable.
Test Time Inference. We employ the teacher model $F_{t}$ at test time. Our method does not require warping or unwarping during testing. This contrast with image warping techniques proposed in the supervised setting [38, 10, 39], which incur overheads from test-time warping and require additional information (like vanishing points [10]) to generate saliency maps for warping at test time.
Supervised Setting. Our training method can also be applied in supervised learning scenario by following the supervised pre-training steps to train a model $F^{\prime}$. That is, omitting the unsupervised training stage completely and only performing supervised pre-training stage – then removing the warping and unwarping layers completely from the trained model. This is unlike prior work [27, 38, 10, 39] which also advocates warping at test time. With our warping on the source domain, model performance improves on the target domain $\mathcal{X}_{b}$ even without adaptation (See Table 5 and 6). For Source Domain Results, see Supp. C.
@@ 4 Experiments
We implement saliency-guided warping alongside a base domain adaptation algorithm, ensuring fair comparisons by using the same model, datasets, training schedules, hyperparameters, and seeds. We compare with 2PCNet [19] for detection, and DAFormer [15], HRDA [16] and MIC [17] for segmentation. More details are provided in Supp. F.
@@ 4.1 Domain Adaptive Object Detection
Tables 123 and 4 show our method improves the state-of-the-art 2PCNet [19] across lighting, weather and geographic adaptations. Besides, our instance-level saliency guidance consistently outperforms approximate saliency guidance [38, 10]. Note that 2PCNet [19] focused solely on Day $\rightarrow$ Night adaptation with its NightAug augmentation, which we only adopted for BDD100K Day $\rightarrow$ Night (Table 3).
Figure 6 shows examples where 2PCNet [19] mistakes parts of the roads as cars and streetlights as traffic lights (For other domain adaptations, see Supp. E). In contrast, our method accurately distinguish foreground objects from background elements in challenging lighting and weather.
@@ 4.2 Domain Adaptive Semantic Segmentation
Our approach is task-agnostic, allowing integration into DAFormer [15], HRDA [16] and MIC [17] for domain adaptive semantic segmentation. We observe significant improvements on ACDC and DarkZurich dataset (Table 5 and Table 6) via integration our method into all these algorithms. Besides, our instance-level image saliency guidance outperforms Static Prior [38] and Geometric Prior [10]. For comparisons with HRDA and MIC, see Supp. B. Note that HRDA and MIC use heavy resize and multi-crop augmentations during training and sliding window inference at test, significantly increasing training memory (+5 GB) and inference latency (+800 ms on an RTX 4090 GPU). In contrast, our approach requires minimal additional training memory (+0.04 GB), negligible training latency (+5.7 ms on an RTX 4090 GPU) and no additional inference latency.
Tables 5 and 6 also demonstrates our method’s superiority across different scenarios (w/ and w/o UDA), indicating that the learnt backbone features generalize better to unseen domains, since the backbone model already improves after supervised pre-training with image warping solely on source. A visual diagnosis of the learned backbones can be found in Supp. E, including GradCam [34] visualizations showing higher discriminability and an increased focus on salient objects when saliency guided warping is employed while training. Additionally, Figure 7 illustrates the superior performance of our method in identifying both background (e.g., sky, sidewalks, buildings) and objects (e.g., traffic lights, traffic signs).
@@ 5 Limitations and Conclusion
This work demonstrates that saliency-guided in-place image warping enhances Unsupervised Domain Adaptation (UDA). Unlike prior approaches that uniformly process objects and backgrounds, our instance-level saliency guidance oversamples object regions and undersamples backgrounds, improving backbone feature learning and aiding domain adaptation. Extensive experiments validate the effectiveness and efficiency of our approach, which adds minimal training overhead, no inference latency, and is agnostic to the adaptation task, target domain, saliency guidance, model architecture, and domain adaptation algorithm.
A limitation of our method is its reduced effectiveness in densely populated scenes where there is limited background to reduce for enlarging foreground objects. Our in-place warping strategy is less effective when background is less than one-fourth of the image. Besides, our approach is less effective on synthetic datasets like Synthia [30]. However, it proves effective on other synthetic datasets like GTA5 [29], which feature realistic city driving scenes.
Acknowledgment: This work was supported by a research contract from General Motors Research-Israel, NSF Grant CNS-2038612, a US DOT grant 69A3551747111 through the Mobility21 UTC and grants 69A3552344811 and 69A3552348316 through the Safety21 UTC.
@@ Appendix A Synthetic to Real Domain Adaptation
Table 9: Synthetic to Real Domain Adaptation: GTA $\rightarrow$ Cityscapes Semantic Segmentation. Tested on Cityscapes Val. We were unable to reproduce DAFormer’s [15] results averaged over 3 random seeds. Therefore, we present results with seed=0 for both DAFormer [15] and our method (DAFormer with instance-level saliency guidance). Our method shows improvement in many individual categories for GTA. Although our focus is on real datasets rather than synthetic datasets, GTA closely resembles natural driving scenes due to its realistic traffic simulation and vehicle-perspective imagery (Figure 8). These realistic object-background appearances with background dominance contribute to the effectiveness of our method.
@@
Table 10: Synthetic to Real Domain Adaptation: Synthia $\rightarrow$ Cityscapes Semantic Segmentation. Tested on Cityscapes Val. We were unable to reproduce DAFormer’s [15] results averaged over three random seeds. Therefore, we present results with seed=0 for both DAFormer [15] and our method (DAFormer with instance-level saliency guidance). Our improvement on Synthia $\rightarrow$ Cityscapes is marginal, which we attribute to the fact that Synthia does not exhibit appearance akin to natural driving scenes (Figure 8). These unrealistic object-background appearances make our method ineffective.
@@ Appendix B Comparison with other Domain Adaptation Strategies
Table 11: Comparison with other Domain Adaptation Strategies: Computational Efficiency. We use a single RTX 4090 GPU with a batch size of 1, maintaining the same training and inference image size of 512x1024 as specified in DAFormer [15]. We also include a comparison using a single A40 GPU, which has more memory, enabling us to experiment with an image size of 768x1536.
Our saliency-guided image warping imposes minimal memory overhead during training. Additionally, since we do not perform warping during test time, our method incurs no inference latency overhead. Image Size Method Training Memory Inference Time 512x1024 DAFormer 16.41 GB 195.0 ms DAFormer + Ours + 0.04 GB + 0 ms DAFormer + HRDA + 5.11 GB + 801.2 ms 768x1536 DAFormer 30.40 GB 261.3 ms DAFormer + Ours + 0.04 GB + 0 ms DAFormer + HRDA + 6.69 GB +1074.0 ms
@@ B.2 Computational Efficiency
Table 11 shows the computational efficiency comparison using the DAFormer image scale. We observe that incorporating HRDA into DAFormer significantly increases training and inference computational costs due to HRDA’s train-time HR-cropping and test-time sliding window inference. In contrast, our inference costs remain the same as those of DAFormer [15], as we do not warp during test time. Our training memory consumption are only slightly higher, due to the lightweight design of our warping-related modules. As we noted in the main manuscript, no additional learned parameters are introduced. Moreover, the latency of the warping and unwarping operations is minimal, at 1.5 ms and 4.2 ms respectively.
@@ B.3 Evaluation Methodology
Based on the above discussion, we apply HRDA’s and MIC’s strategies to DAFormer. However, for a fair comparison, we perform training and inference following DAFormer’s [15] training/testing image scales and evaluation paradigm. An alternative comparison would involve training and evaluating on full-scale images to match HRDA, which deviates from original DAFormer [15] training and evaluation that uses half-scaled images. Unfortunately, we lack access to a GPU with high memory capacity, such as Tesla A100 with 80 GB of memory.
@@ Appendix E Additional Analysis
Appendix F Experimental Methodology
Semantic Segmentation: We follow DAFormer [15] by employing the same SegFormer [49] head and MiT-B5 [49] backbone, adhereing to their training hyperparameters, protocols, and seed for fair comparison. While Sim2Real Gap result presented in DAFormer [15] are not our focus, relevant results and discussion can be found in Supp. A.
Datasets: We use BDD100K [52], Cityscapes [6], DENSE [3], ACDC [33] & DarkZurich [31]. A brief description is given below.